# Optimization under Attack: Resilience, Vulnerability, and the Path to Collapse

### v1

@@ Abstract
Optimization is instrumental for improving operations of large-scale socio-technical infrastructures of Smart Cities, for instance, energy and traffic systems. In particular, understanding the performance of multi-agent discrete-choice combinatorial optimization under distributed adversary attacks is a compelling and underexplored problem, since multi-agent systems exhibit a large number of remote control variables that can influence in an unprecedented way the cost-effectiveness of distributed optimization heuristics. This paper unravels for the first time the trajectories of distributed optimization from resilience to vulnerability, and finally to collapse under varying adversary influence. Using real-world data to emulate over 28 billion multi-agent optimization scenarios, we exhaustively assess how the number of agents with different adversarial severity and network positioning influences optimization performance, including the influence on Pareto optimal points. With this novel large-scale dataset, made openly available as a benchmark, we disentangle how optimization remains resilient to adversaries and which adversary conditions are required to make optimization vulnerable or collapsed. These new findings can provide new insights for designing self-healing strategies for fault-tolerance and fault-correction in adversarial distributed optimization that have been missing so far.
@@ I Introduction
The rapid development of Internet of Things (IoT) applications has brought about transformative changes in numerous domains, ranging from smart cities to industrial automation and healthcare [1]. These applications involve vast networks of interconnected devices that generate substantial amounts of data, and require efficient and decentralized decision-making for processing [2]. Distributed optimization is essential in such scenarios, as it enables multiple agents to collaborate effectively without centralized coordination. This approach ensures scalability, reliability, and robust performance across diverse applications, including energy management, autonomous systems, and federated learning [3, 4, 5, 6, 7].
@@
While distributed optimization is essential for achieving system-wide objectives, most algorithms rely on assumptions of rational, cooperative, and non-adversarial agents contributing toward the global objective without prioritizing their individual goals over collective outcomes. However, real-world scenarios often deviate from these assumptions, as adversarial agents can disrupt optimization systems by introducing inaccuracies, manipulating decisions, or compromising functionality [8, 9, 10, 11]. For example, in smart grid systems, where autonomous agents (e.g., households) collaborate to manage energy distribution and prevent power outages, certain households may act adversarially by manipulating consumption data and prioritizing self-interest, i.e. their thermal comfort. Such disruptions distort energy allocation processes, leading to inefficient resource use and blackouts [9, 10, 12, 13, 14].
@@
Several studies have examined multi-agent optimization in continuous-choice frameworks [8, 9, 10, 11, 13, 15]. While stochastic optimization techniques address noise, they lack mechanisms to counteract strategic manipulations by adversarial agents [16, 17]. These manipulations amplify system vulnerability by prioritizing individual goals over collective objectives, which can lead to inefficiencies, instability, and eventual collapse [18, 19]. Despite this, adversarial disruptions in discrete-choice settings have received little attention.
@@
This paper addresses this gap by investigating the resilience, vulnerability, and collapse of multi-agent systems in discrete-choice optimization under adversarial conditions, offering a novel framework for understanding system behavior. Resilience refers to the system ability to maintain performance despite adversarial influence. Vulnerability captures the emergence of inefficiencies due to adversarial behaviors, and collapse occurs when the system fails to achieve its objectives. This study evaluates the impact of adversarial agents on optimization performance, analyzing critical parameters such as the number of adversarial agents, their adversarial severity, and their network positions. By systematically examining these dynamics, the paper provides critical insights into the conditions under which systems transition from resilience to collapse, unravelling new insights for the development of strategies [20] for self-healing, fault-tolerance, and fault-correction in adversarial environments.
@@
  • A novel evaluation framework for characterizing resilience, vulnerability, and collapse in distributed optimization systems under adversarial conditions.
  • The applicability of the proposed adversarial model in different application domains.
  • Open large-scale benchmark datasets comprising over 13 million experiments on real-world data, for comprehensive evaluation of distributed optimization systems in the presence of adversaries.
@@
  • New insights highlighting adversarial influences on optimality in various application scenarios.
  • An open-source software artifact implementation of the proposed adversarial model for the I-EPOS collective learning algorithm [^1] Available at https://github.com/epournaras/EPOS.
The rest of this paper is organized as follows: Section II reviews related literature on adversarial impacts in distributed optimization. Section III introduces the proposed adversarial model, along with the problem formulation, detailing the network model and optimization challenge. Section IV outlines the experimental methodology and evaluation metrics. The key findings are presented in Section V, discussing system behavior under varying adversarial conditions. Finally, Section VI concludes the paper and outlines directions for future research.
@@ II Related Work
Resilience in distributed multi-objective optimization plays a critical role across domains such as smart grids, transportation, logistics, and communication networks, where robust and adaptive systems are crucial for ensuring operational efficiency [21, 22]. Convex distributed optimization has received significant attention, with a focus on addressing challenges posed by adversary agents, network structures, and varied application domains [3, 13, 23]. Earlier work examines the robustness and vulnerability of consensus-based distributed optimization to address limitations and assumptions related to adversary behavior, network topology, objective functions, and application domains [13, 24, 25]. The presence of adversary agents significantly impacts the performance of distributed optimization models. These agents disrupt optimization by slowing convergence, manipulating data, or withholding participation, resulting in suboptimal performance [8, 15]. Table I provides a comparative analysis of related work in distributed optimization under adversarial conditions, focusing on adversary behavior, attack strategies, and their impact on network and system performance.
@@
Work Adversarial behavior Adversary target Knowledge of the system Network Performance measure Algorithms / Techniques for Optimization Recovery [10] Malicious & Byzantine  [^2] A node is considered malicious if it sends the same value to all its out-neighbors at each time-step, while a Byzantine node can send different values to different neighbors at each time-step. Consensus protocol Full knowledge of network structure and cost functions of other agents Connected undirected network Distance to optimality minimization Local filtering [26] Byzantine Consensus process & parameter agreement Full knowledge of network structure and cost functions of other agents Strongly connected directed network Convergence to a bounded region containing the global minimizer Distance-based filtering [27] Byzantine Information exchange among agents Full knowledge of network structure and cost functions of other agents Complete directed network Admissibility (structure of the convex coefficients) Local filtering [28] Adversarial manipulation of observations Single agent observation Black-box access to an agent’s policy and environment Limited directed communication Average reward & team win rate Gradient-based methods (Deep Q-learning) [29] Malicious Consensus-based MARL Partial knowledge (compromises rewards) Fully connected network Average team reward Consensus-based MARL [30] Data perturbations (cyber-attacks) DRL rewards Full knowledge of DRL model parameters DNN (power system topology) Expected performance decay & critical attack rate None [31] Injection & DoS attacks Consensus in cyber-physical systems Full knowledge of network structure and cost functions of neighbors Robust sparse network Resilient asymptotic consensus Mean subsequence reduced [32] Malicious data injections Convergence to the global optimal point Partial knowledge of network structure and neighbors’ states Undirected connected network Convergence to optimal point Probabilistic trust-based weight adjustment & projection-based update [33] Malicious & Byzantine Information exchange among agents Full knowledge of network structure and neighbors’ states Directed connected spanning network Convergence to optimal point Markov switching communication topology & Push-DIGing [34] Malicious & Byzantine Consensus process and state updates Full knowledge of network structure and neighbors’ states Undirected connected network Convergence to optimal point Resilient distributed optimization with trusted agents & connected dominating set [35, 36] Impersonation Byzantine Communication channels between agents and coordinator Local knowledge Directed network with central coordinator Distance to optimality minimization Primal-Dual [9] Malicious Consensus value Local knowledge Directed & strongly connected Distance to optimality minimization FROST [37] Malicious Consensus value Local knowledge Undirected network Distance to optimality minimization Randomized gossip [38, 39] Eavesdropping Information exchange among agents All transmitted messages Undirected network Convergence to a privacy-preserving approximate solution Differentially private gradient tracking This work Selfish malicious System-wide optimization Local knowledge & black-box access to network structure Bidirectional hierarchical network Optimization efficiency I-EPOS
@@ II-B Adversarial Attacks in Multi-Agent Systems
Adversarial attacks significantly impact reinforcement learning (RL) systems across applications such as robotics, video games, and smart grids, undermining system stability and performance [40, 41]. Lin et al. [28] demonstrate how adversarial perturbations affect cooperative multi-agent RL (c-MARL), showing its vulnerability compared to single-agent RL. Figura et al. [29] highlight how a single adversary can influence consensus-based c-MARL systems, disrupting team objectives. Zheng et al. [30] introduce criticality-based perturbations in deep Q-networks, demonstrating substantial performance degradation due to adversarial attacks. These studies highlight that an adversary agent has the ability not only to disrupt system operations but also to manipulate the policy to align with its own adversarial objectives, thereby influencing other agents to adopt a policy of its preference. These insights underscore the critical need for robust defense against adversaries that manipulate policies and system states in RL systems.
@@ II-D Resource Allocation Challenges Under Adversaries
In resource allocation, adversarial disruptions are mitigated using robust optimization algorithms. Uribe et al. [35] and Turan et al. [36] propose primal-dual methods to tolerate adversaries by identifying and eliminating Byzantine inputs. Similarly, Ravi et al. [9] and Gentz et al. [37] focus on adversary detection, using statistical analysis and hypothesis testing to isolate malicious agents. These methods achieve resilience for up to 50% adversary density. However, they assume adversarial influence is evenly distributed and may not generalize to more concentrated or dynamic adversarial settings.
Another approach focuses on devising an adversary detection model to identify malicious agents in distributed optimization systems [9, 37]. Ravi et al. [9] address the challenge of detecting and mitigating malicious inputs in a distributed optimization system. Their method focuses on utilizing the data values of the agents to identify and isolate potential malicious agents. The researchers impose an upper limit on the number of tolerable malicious agents, specifically setting it at half of the total number of agents in the network topology. On the other hand, Gentz et al. propose a detection method for insider attackers in randomized gossip-based sensor networks. This approach involves utilizing theory and hypothesis testing to analyze the statistical states of the sensors, enabling the detection of malicious agents within the network [37].
II-E Privacy-Preserving Distributed Optimization
@@
Privacy-preserving distributed optimization safeguards sensitive agent information against eavesdropping adversaries using techniques such as differential privacy [38, 39], homomorphic cryptography [42, 43], and gradient perturbation [44, 45], to ensure secure information exchange. However, these studies focus on privacy protection rather than adversarial behavior in optimization contexts.
II-F Combinatorial Optimization Algorithms
@@
In distributed collective learning and combinatorial optimization, Hinrichs et al. [46, 47] propose COHDA, a combinatorial optimization heuristic designed for multi-agent systems. However, COHDA encounters scalability challenges due to increasing communication overhead as the number of agents grows. Pournaras et al. [48] address this with EPOS, a decentralized framework that enables agents to collaboratively optimize global resource allocation, particularly in participatory sharing economies. EPOS ensures privacy, autonomy, and scalability but encounters computational limitations when applied to wide tree structures with multiple child nodes [49]. I-EPOS extends EPOS by introducing decentralized iterative back-propagation and localized decision-making to enhance scalability [50, 49]. While these works significantly contribute to decentralized combinatorial optimization, none of them study the performance in the presence of adversary agents.
Research on adversarial attacks predominantly focuses on continuous distributed optimization, often analyzing specific attack types or assuming a limited number of adversaries, which may not accurately reflect real-world scenarios. These studies typically rely on complete graph assumptions and address single points of failure rather than system-wide resilience. While several solutions mitigate adversarial attacks in multi-agent optimization, there is a notable lack of comprehensive analyses of inherent system vulnerability, resilience, and how optimization collapses.
This study addresses the critical gap in understanding how adversary agents impact system functionality by introducing a generic adversarial model to analyze resilience and vulnerability in distributed multi-objective optimization. The primary objective is to assess in-depth the influence of adversary agents on system optimality. By filling this gap, this research aims to advance the design of self-healing strategies to enhance fault-tolerance and fault-correction in adversarial distributed optimization, which, to the best of our knowledge, has not been comprehensively explored in prior research.
@@ III-A Problem Formulation
Resilience in distributed optimization is essential for maintaining system performance under adversarial conditions. Adversary agents disrupt operations, degrade efficiency, and increase system vulnerability. This raises key questions: How do adversary agents influence the efficiency and stability of distributed optimization systems? What thresholds of adversarial behavior lead to transitions from resilience to vulnerability or collapse? How do parameters such as adversary density, adversarial severity, and network positioning influence these transitions?
To address these challenges, we propose a generic adversarial distributed optimization model tailored to discrete-choice scenarios. This model is used to investigate the trade-offs between system-wide and individual agent goals in adversarial settings. It incorporates critical parameters, including adversary density, adversarial severity, and structural positioning, providing a robust framework to evaluate system vulnerability and resilience. This study identifies thresholds for transitions from resilience to collapse, providing an in-depth understanding of system behavior under adversarial influence. These findings inform the development of self-healing strategies that enhance fault-tolerance and mitigate adversarial impacts across diverse distributed optimization applications.

### v2

@@ Abstract
Optimization is critical for improving the operations of large-scale socio-technical infrastructures such as those found in energy, mobility, and information systems. In particular, understanding the performance of multi-agent discrete-choice combinatorial optimization under distributed adversarial attacks is a compelling and underexplored problem. Multi-agent systems involve a large number of remote control variables that can influence the cost-effectiveness of distributed optimization heuristics. This paper unravels, for the first time, the trajectories of distributed optimization from resilience to vulnerability, and finally to collapse under varying adversarial influence. Using real-world and synthetic data to generate over 112 million multi-agent optimization scenarios, we systematically assess how the number of agents with varying levels of adversarial severity and network positioning influences optimization performance, with particular attention to the impact on Pareto optimality. With this large-scale dataset, made openly available as a benchmark, we disentangle how optimization systems remain resilient to adversaries and which adversary conditions make optimization vulnerable or cause collapse. These findings can support the design of self-healing strategies for fault tolerance and fault correction, addressing a critical gap in adversarial distributed optimization.
@@ I Introduction
The rapid development of Internet of Things (IoT) applications has brought transformative changes in numerous domains, ranging from smart cities to industrial automation and healthcare [1]. These applications involve vast networks of interconnected devices that generate substantial amounts of data, and require efficient and distributed decision-making [2]. Distributed optimization is essential in such scenarios, as it enables multiple agents to collaborate effectively without centralized coordination. This approach ensures scalability, reliability, and robust performance across diverse applications, including energy management, autonomous systems, and federated learning [3, 4, 5, 6, 7].
@@
While distributed optimization is essential for achieving system-wide objectives, most algorithms rely on assumptions of rational, cooperative, and non-adversarial agents contributing toward the global objective without prioritizing their individual goals over collective outcomes. However, real-world scenarios often deviate from these assumptions; adversary agents can disrupt optimization by introducing inaccuracies, manipulating decisions, or compromising functionality [8, 9, 10, 11]. For example, in smart grid systems where autonomous agents (e.g., households) collaborate to manage energy distribution and prevent power outages, certain households may act adversarially by manipulating consumption data to prioritize their self-interest, i.e., their thermal comfort. Such disruptions distort energy allocation, leading to inefficient resource use and blackouts [9, 10, 12, 13, 14].
@@
Several studies have examined multi-agent optimization in continuous-choice frameworks [8, 9, 10, 11, 13, 15]. While stochastic optimization techniques address noise, they lack mechanisms to counteract strategic manipulations by adversarial agents [16, 17]. Such manipulations amplify system vulnerability by prioritizing individual goals over collective objectives, which can lead to inefficiencies, instability, and eventual system collapse [18, 19]. Despite these risks, adversarial disruptions in discrete-choice settings have received little attention.
@@
This paper studies multi-agent systems in discrete-choice combinatorial optimization under adversarial conditions, with the aim to unravel the trajectories of resilience, vulnerability, and collapse, offering a novel framework to understand system optimization behavior. In this context, resilience refers to the system ability to maintain performance despite adversarial influence; vulnerability captures the emergence of inefficiencies due to intolerable adversarial behaviors; and collapse occurs when the system significantly under-performs as a result of failure to cope with adversarial behaviors. This study evaluates the impact of adversarial agents on optimization performance, analyzing critical parameters such as the number of adversarial agents, their behavioral severity, and their network positions. By systematically examining these dynamics, the paper provides critical insights into the conditions under which systems transition from resilience to collapse, offering new insights for developing self-healing, fault-tolerant, and fault-correcting strategies in adversarial environments [20].
@@
  • A novel evaluation framework for characterizing resilience, vulnerability, and collapse in distributed optimization systems under adversarial influence.
@@
  • The first open large-scale benchmark datasets for discrete-choice adversarial optimization, generated from over 112 million experiments on real-world and synthetic inputs using the proposed adversarial model.
  • New insights into how adversarial scale, severity, and structural positioning affect system optimality, including resilience and vulnerability thresholds, Pareto trade-offs, and structural vulnerabilities across diverse scenarios.
  • An open-source software artifact implementing the proposed adversarial model within the I-EPOS system [^1] Available at https://github.com/epournaras/EPOS, extending its functionality to support heterogeneous agent behaviors.
The rest of this paper is organized as follows: Section II reviews related work on adversarial distributed optimization. Section III introduces the proposed adversarial model and problem formulation, including the network model and optimization challenge. Section IV outlines the experimental methodology and evaluation metrics. Section V presents the key findings, analyzing system behavior under varying adversarial conditions. Finally, Section VI concludes the paper, discusses its limitations, and outlines directions for future research.
@@ II Related Work
Resilience in distributed multi-objective optimization plays a critical role across domains such as smart grids, transportation, logistics, and communication networks, where robust and adaptive systems are crucial for ensuring operational efficiency [21, 22]. Convex distributed optimization has received significant attention, with a focus on addressing challenges posed by adversary agents, network structures, and varied application domains [3, 13, 23]. Earlier work examined the robustness and vulnerability of consensus-based distributed optimization, focusing on addressing limitations related to adversary behavior, network topology, objective functions, and application domains [13, 24, 25]. The presence of adversary agents significantly impacts the performance of distributed optimization models. These agents disrupt optimization by slowing convergence, manipulating data, or withholding participation, resulting in suboptimal performance [8, 15]. Table I provides a comparative analysis of related work on distributed optimization under adversarial conditions. It highlights key aspects such as the type of adversarial behavior [^2] A malicious node sends the same value to all its neighbors at each time step, whereas a Byzantine node may send different values to different neighbors., attack targets, system knowledge [^3] Knowledge levels: full—complete knowledge of the network and agent objectives; partial—access to limited neighbor information; local—only own state or data is known., network structures, and the impact on overall performance.
@@
[10] [26] [27] [28] [29] [30] [31] [32] [33] [34] [35, 36] [9] [37] [38, 39] This work Adversarial behavior Malicious Byzantine Cyber‐attacks Eavesdropping Adversary target Consensus Information Exchange System Objective Observation Knowledge of the system Full Partial Local Directed Network Performance measure Convergence Distance to Optimality Reward/Utility Efficiency Algorithms / Techniques for Optimization Local filtering Distance– based filtering Local filtering Gradient –based (Deep Q‐learning) Consensus‐based MARL None Mean subsequence reduced Probabilistic trust– based & projection–based Markov switching topology & Push‐ DIGing Resilience with trusted agents & dominating set Primal– Dual FROST Randomized gossip Differentially private gradient tracking I‐EPOS
@@ II-B Adversarial Attacks in Multi-Agent Systems
Adversarial attacks significantly impact reinforcement learning (RL) systems across applications such as robotics, video games, and smart grids, undermining system stability and performance [40, 41]. Lin et al. [28] demonstrate how adversarial perturbations affect cooperative multi-agent RL (c-MARL), showing its vulnerability compared to single-agent RL. Figura et al. [29] highlight how a single adversary can influence consensus-based c-MARL systems, disrupting team objectives. Zheng et al. [30] introduce criticality-based perturbations in deep Q-networks, demonstrating substantial performance degradation due to adversarial attacks. These studies show that an adversary can disrupt system operations and manipulate policies, influencing other agents to adopt behaviors aligned with its objectives.
@@ II-D Resource Allocation Challenges Under Adversaries
In distributed resource allocation, adversarial disruptions are typically mitigated using robust optimization and detection mechanisms. Uribe et al. [35] and Turan et al. [36] propose primal-dual methods that tolerate Byzantine adversaries by identifying and eliminating malicious inputs, achieving resilience for up to 50% adversary density. Similarly, Ravi et al. [9] develop a detection method that uses agents’ data values to identify and isolate potential malicious behavior, imposing an upper limit of 50% adversaries within the network topology. Gentz et al. [37] propose a detection method based on hypothesis-testing for insider attackers in randomized gossip-based sensor networks, leveraging statistical analysis of sensor states to identify malicious agents. While these methods enhance resilience against dispersed adversaries, they assume adversarial influence is evenly distributed and may not generalize to scenarios with concentrated or dynamic adversary placement.
II-E Multi-Objective Distributed Optimization
Recent studies have increasingly adopted Pareto-based multi-objective optimization to evaluate system trade-offs in complex infrastructures. Fettah et al. [42] introduce a Pareto strategy for optimizing distributed generation in power networks. Zhang et al. [43] formulate a multi-objective operational framework that integrates Pareto analysis to enhance resilience thresholds in distribution networks. Similarly, Boindala and Ostfeld [44] propose an optimization approach to balance reliability, cost, and failure risk using Pareto fronts. While these studies underscore the value of Pareto analysis for resilient optimization, they focus on centralized infrastructures and do not address adversarial influence or the complexities of distributed, multi-agent decision-making explored in this work.
II-F Privacy-Preserving Distributed Optimization
@@
Privacy-preserving distributed optimization safeguards sensitive agent information against eavesdropping adversaries using techniques such as differential privacy [45, 38, 39], homomorphic cryptography [46, 47], and gradient perturbation [48, 49], to ensure secure information exchange. However, these studies focus on privacy protection rather than adversarial behavior in optimization contexts.
II-G Combinatorial Optimization Algorithms
@@
In distributed combinatorial optimization, Hinrichs et al. [50, 51] propose COHDA, a combinatorial optimization heuristic designed for multi-agent systems. However, COHDA encounters scalability challenges due to increasing communication overhead as the number of agents grows. The collective learning approach of Pournaras et al. [52] address this with EPOS (Economic Planning and Optimized Selections), a distributed optimization method that enables agents to collaboratively optimize global resource allocation, particularly in participatory sharing economies. Although EPOS ensures privacy, autonomy, and scalability, it faces computational limitations when applied to wide tree structures with multiple child nodes [53]. I-EPOS is the iterative extension of EPOS; it incorporates decentralized iterative back-propagation and localized decision-making to enhance scalability and support plan coordination across deeper and broader network hierarchies [54, 53].
While COHDA, EPOS, and I-EPOS represent foundational combinatorial optimization approaches, their system performance under adversarial conditions has not been studied before, despite some limited work on measuring the impact of arbitrary structural faults, which does not focus on agents’ behavior [55]. In this work, we introduce a generic adversarial model applicable to such settings. The model enables a structured evaluation of resilience, vulnerability, and collapse dynamics, and, to the best of our knowledge, is the first to systematically explore adversarial behaviors in discrete-choice combinatorial optimization.
A large body of research to date focuses on continuous distributed optimization, often assuming limited adversary presence or relying on complete graph topologies [8, 9, 10, 13]. Such assumptions do not fully capture the complexity of real-world systems, where distributed structures, heterogeneous agent behaviors, and dynamic adversarial threats are prevalent [30, 10]. Although existing solutions offer valuable mitigation strategies [35, 33], the lack of comprehensive analyses on inherent system vulnerability, resilience thresholds, and pathways to optimization collapse remains a gap.
To fill this gap, we propose a generic adversarial model to systematically analyze how adversarial agents influence system performance and stability in distributed multi-objective optimization.Our objective is to evaluate how adversarial behavior influences resilience, vulnerability, and collapse, while informing the development of self-healing strategies for robust optimization. Moreover, his work releases the first large-scale, open benchmark dataset designed for evaluating adversarial impacts under discrete-choice optimization settings. To clarify the novelty of our contribution, Table II summarizes a comparative analysis with existing work, highlighting key aspects such as discrete decision-making, multi-objective formulation, resilience thresholds, structural vulnerability, and Pareto-based evaluation.
TABLE II: Comparison of the novelty aspects of this work with related distributed optimization approaches
Criteria This work [10] [26] [27] [28] [29] [30] [31] [32, 33, 34] [35, 36] [9] [37] [38, 39] [42, 43, 44] Multi-objective optimization X X X X X X Discrete decision-making X X X X X X X X X X Resilience & vulnerability thresholds X X X X Structure analysis X X X X X X X X X Pareto analysis X X X X X X X X X X X X
@@ III-A Problem Formulation
Resilience in distributed optimization is essential for maintaining system performance under adversarial conditions. Adversary agents disrupt operations, degrade efficiency, and increase vulnerability. This raises key questions: How do adversary agents influence the efficiency and stability of distributed optimization systems? What thresholds of adversarial behavior lead to transitions from resilience to vulnerability or collapse? How do parameters such as adversary density, adversarial severity, and network positioning influence these transitions?
To address these challenges, we propose a generic adversarial distributed optimization model tailored to discrete-choice scenarios. Our model investigates the trade-offs between system-wide and individual agent goals in adversarial settings. It incorporates key parameters, including the scale of adversaries, behavioral severity, and structural positioning, providing a robust framework to evaluate system vulnerability and resilience. Through this study, we identify critical thresholds for transitions from resilience to collapse, providing an in-depth understanding of system behavior under adversarial influence. These findings inform the development of self-healing strategies that enhance fault-tolerance and mitigate adversarial impacts across diverse distributed optimization applications.